# Center vortices and localized Dirac modes in the deconfined phase of 2+1 dimensional lattice $\mathbb{Z}_{2}$ gauge theory

### v1

@@ III.4 Refined sea/islands picture
The correlation between islands and modes is studied by looking at ${\cal A}_{\psi}$, i.e., ${\cal A}(\vec{x})$ “as seen by the modes” normalized by $\langle{\cal A}\rangle$, Eq. (30). This is shown in Fig. 17. For both low and high modes, ${\cal A}_{\psi}$ is larger than 1 in both phases. In the deconfined phase the ratio shoots up for low modes, showing that they tend to localize where ${\cal A}(\vec{x})$ is larger. High modes are localized already in the confined phase, but for them the ratio shoots up in the deconfined phase even more than for low modes. For bulk modes instead ${\cal A}_{\psi}$ is smaller than 1 in both phases, indicating that these modes are repelled by the islands.
@@
where $c_{n}(t,\vec{x})=1$ if the site $(t,\vec{x})$ on the direct lattice is a corner of a plaquette whose corresponding dual link belongs to cluster $n$, and $n_{c}(t,\vec{x})$ is the number of clusters that touch site $(t,\vec{x})$. For the correlation between islands and Polyakov loops we used $C_{\mathcal{A}P}$ defined in Eq. (31), measured on 700 configurations. Both correlations become very strong in the deconfined phase. [^3] For a quantitative comparison between the magnitudes of the two types of correlation one should normalize by the fluctuations of the various quantities instead of their averages. On the other hand, an accurate quantitative comparison is not really possible: while for $C_{\mathcal{A}P}$ both quantities are defined on the direct lattice sites, $O_{C\mathcal{A}}$ involves a quantity defined on the direct lattice and one defined on the dual lattice. This strongly suggests that the “islands” relevant to localization are found at the location of those center vortices generated by Polyakov-loop fluctuations away from order.
@@ III.5 Center vortex structures and localization
Figure 20: Comparison of the normalized spectral density $\rho(\lambda)$, Eq. (10), at $\beta=0.74$ on a $72^{2}\times 4$ lattice with the localized low modes of the simplest vortex configurations of Tab. 1.
@@
Here objects are counted by identifying exact matches between disconnected components of the vortex configuration on the dual lattice and the configurations shown in Tab. 1, possibly rotated or reflected in space, and reflected in the temporal direction. For a given type of object, a strong correlation is found precisely at the position of the corresponding localized low and high eigenmodes, $\lambda_{O}$ and $\tilde{\lambda}_{O}$, found in the presence of a single, isolated object; a clear anticorrelation is also seen near the mobility edges. [^5] We did not determine precisely the position of the mobility edge separating delocalized bulk and localized high modes, but all observables point at it being found near the high end of the bulk. The position of the peaks in the correlation match almost perfectly those in the spectral density of low modes, as shown in Fig. 21 (second panel), while the matching seems less nice at the high end of the spectrum (Fig. 21, third panel). However, although relatively rare, the objects are not isolated from each other, and so the spectrum is not simply the union of the individual spectra of the configurations corresponding to the objects in Tab. 1, and the positions of the peaks in $\rho$ are expected to deviate from the isolated-object case. On the other hand, a comparison between the positions of the correlation peaks and the eigenvalues of the localized modes for isolated objects (fourth panel) shows that even in a heterogeneous environment the spectrum responds the most to the center-vortex configurations of Tab. 1 almost exactly where the “free” eigenvalue would be. The bigger deviation of the spectral density peaks from the correlation peaks for the high modes is then likely due to the fact that the localized high modes associated with isolated objects have eigenvalues closer to each other than the corresponding localized low modes, and therefore are more likely to mix when several objects are present, distorting the peak structure from the isolated-object case more than at the low end of the spectrum.
@@
Figure 22: Normalized spectral density at fixed number $\mathcal{N}$ of “vertical squares” (top panel), and its difference with the full normalized spectral density (bottom panel). The dashed line is at $\lambda\simeq 0.55$, which is the eigenvalue of the localized low mode corresponding to an isolated vertical square; the solid line indicates the position of the mobility edge. Here $\beta=0.74$, $N_{s}=20$, and $N_{t}=4$.
@@
Since $N_{O}V_{O}/V<1$ is the fraction of the lattice occupied by objects, the correlation is positive. As the temperature grows this fraction decreases, and so does the mode size (see Fig. 1), approximated here by $n_{O}V_{O}$, leading therefore to stronger correlations.
Interestingly, in Fig. 24 a positive correlation between low and high modes, particularly strong near $(\lambda_{1},\lambda_{2})=(\lambda_{O},\tilde{\lambda}_{O})$, is observed also in the confined phase. As a nonzero density of finite clusters is present also in the confined phase (see Sec. III.2), it is not surprising that they support the nonzero spectral density of localized high modes, as this result further confirms. It is somewhat surprising, though, that a strong correlation with the spectral regions where the low-mode counterparts would be found survives also in the confined phase, where low modes are delocalized. This indicates that low modes still carry some local structure highly correlated with the center-vortex configuration. This also explain their anticorrelation with the bulk modes.
@@
The results of this Subsection show that the localization of low Dirac modes in the deconfined phase of $\mathbb{Z}_{2}$ is explained by the opening of the spectral pseudogap due to Polyakov loop ordering, and by the presence of disconnected finite clusters, i.e., center vortices not spanning the whole lattice, that are the main localization centers. As both Polyakov loop ordering and the fragmentation of the infinite center cluster are direct manifestations of deconfinement, the strong connection between deconfinement and localization is evident. Notice that center vortices being the main localization centers is not at all in contradiction with the sea/islands picture, as Polyakov loop, islands, and center vortices are very strongly correlated with each other, as shown in the previous Subsection.
@@ IV Conclusions
In this paper we investigated different aspects of the deconfinement transition in finite-temperature 2+1-dimensional $\mathbb{Z}_{2}$ gauge theory on the lattice, and of its relation with the localization of low Dirac modes.
As a first step, we confirmed the presence of an Anderson transition in the spectrum of the staggered Dirac operator, separating localized low-lying modes from delocalized bulk modes. The presence of localized low modes, appearing at the deconfinement transition, was shown in Ref. [19], short only of a finite-size scaling analysis. Here we performed this analysis, fully validating the claims of Ref. [19].
Our results are interesting also from the point of view of disordered systems, independently of their connection with the confinement of quarks. Our finite-size-scaling analysis shows that 2+1-dimensional $\mathbb{Z}_{2}$ gauge theory is another (spatially) two-dimensional disordered system in the orthogonal class displaying a BKT-type Anderson transition [50, 52]. Furthermore, the dependence of the position of the mobility edge on the temperature is rather counterintuitive, and different from what has been observed so far in lattice gauge theories.
We then further established the characterization of the deconfinement transition as a percolation transition of center vortices [39, 40], determining the percolation temperature and finding it in agreement with the deconfinement temperature [34]. An interesting finding is that the critical properties of the vortex percolation transition are different from those of ordinary two-dimensional percolation. This is probably due to the long-range nature of the constraint on the percolating objects, i.e., the requirement that vortices form closed loops on the dual lattice.
@@
In 2+1-dimensional $\mathbb{Z}_{2}$ gauge theory the very close connection between confinement and specific topological objects (center vortices, in this case) is almost trivially expected, due to the simplicity of the dynamics (it is nonetheless nice to be able to confirm it explicitly). This allowed us to investigate in an inequivocal way (albeit, perhaps, in too simplified an environment) the connection between deconfinement and localization of Dirac modes, strongly suggested by a wealth of results [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. Specifically, we looked directly at the correlation between localized modes and the center vortices whose percolation signals confinement. Such a correlation is demonstrated here at the microscopic level, identifying the structures in the gauge configurations mainly responsible for localizing the Dirac modes: these are precisely the most elementary center vortices. The resulting picture is then that as the temperature is lowered and the confinement transition is approached, such vortices approach a percolation transition, and the localized modes they support tend to mix more strongly, therefore tending to delocalize. The connection between the confinement/deconfinement transition and delocalization/localization transition is then fully elucidated in terms of the gauge structures responsible for both. This supports our hope that identifying the structures responsible for low-mode localization in more realistic gauge theories, including the physically relevant case of QCD, can lead to a better understanding of the microscopic mechanism behind confinement.
@@ Appendix A Van Hove singularities
Due to the ordering of the Polyakov loop, in the deconfined phase the bulk of the spectrum should resemble that of the free theory in the physical sector, and that of the free theory supplemented by a uniform non-trivial Polyakov loop in the unphysical sector. This includes reproducing the Van Hove singularities [58] of the spectral density where the “band energies” $\mathcal{E}_{n}(\vec{p}\,)=\sqrt{(\sin\omega_{n})^{2}+(\sin p_{1})^{2}+(\sin p_{2})^{2}}$ have a vanishing gradient, $\vec{\nabla}\mathcal{E}_{n}(\vec{p}\,)=[2\mathcal{E}_{n}(\vec{p}\,)]^{-1}(\sin 2p_{1},\sin 2p_{2})=0$. Here the various bands are characterized by $\omega_{n}^{\mathrm{phys}}=\frac{(2n+1)\pi}{N_{t}}$ for the physical sector, and $\omega_{n}^{\mathrm{unphys}}=\frac{2n\pi}{N_{t}}$ for the unphysical sector, with $n=0,\ldots,N_{t}-1$ in both cases. The spatial momentum $\vec{p}=(p_{1},p_{2})$ is restricted to the first Brillouin zone; in the large-volume limit, $p_{i}\in[-\pi,\pi]$.

### v2

@@ III.4 Refined sea/islands picture
The correlation between islands and modes is studied by looking at ${\cal A}_{\psi}$, i.e., ${\cal A}(\vec{x})$ as seen by the modes normalized by $\langle{\cal A}\rangle$, Eq. (30). This is shown in Fig. 17. For both low and high modes, ${\cal A}_{\psi}$ is larger than 1 in both phases. In the deconfined phase the ratio shoots up for low modes, showing that they tend to localize where ${\cal A}(\vec{x})$ is larger. High modes are localized already in the confined phase, but for them the ratio shoots up in the deconfined phase even more than for low modes. For bulk modes instead ${\cal A}_{\psi}$ is smaller than 1 in both phases, indicating that these modes are repelled by the islands.
@@
where $c_{n}(t,\vec{x})=1$ if the site $(t,\vec{x})$ on the direct lattice is a corner of a plaquette whose corresponding dual link belongs to cluster $n$, and $n_{c}(t,\vec{x})$ is the number of clusters that touch site $(t,\vec{x})$. For the correlation between islands and Polyakov loops we used $C_{\mathcal{A}P}$ defined in Eq. (31), measured on 700 configurations. Both correlations become very strong in the deconfined phase. [^4] For a quantitative comparison between the magnitudes of the two types of correlation one should normalize by the fluctuations of the various quantities instead of their averages. On the other hand, an accurate quantitative comparison is not really possible: while for $C_{\mathcal{A}P}$ both quantities are defined on the direct lattice sites, $O_{C\mathcal{A}}$ involves a quantity defined on the direct lattice and one defined on the dual lattice. This strongly suggests that the islands relevant to localization are mainly found at the location of those center vortices associated with Polyakov-loop fluctuations away from order. Notice, however, that only about half of the mode weight is found on the lattice sites corresponding to these fluctuations for the lowest modes, a fraction that decreases as one approaches the mobility edge [19]. This should be compared with the larger fraction found on center vortices, that therefore identify more accurately the localization region of low modes. Moreover and more importantly, as we show in the next subsection, low modes localize mostly where the gauge configuration shows a more complex structure on the dual lattice than for a simple Polyakov-loop fluctuation, indicating that center vortices play a more fundamental role.
@@ III.5 Center vortex structures and localization
Figure 20: Comparison of the normalized spectral density $\rho(\lambda)$, Eq. (10), at $\beta=0.74$ on a $4\times 72^{2}$ lattice with the localized low modes of the simplest vortex configurations of Tab. 1.
@@
Here objects are counted by identifying exact matches between disconnected components of the vortex configuration on the dual lattice and the configurations shown in Tab. 1, possibly rotated or reflected in space, or reflected in the temporal direction, or both. For a given type of object, a strong correlation is found precisely at the position of the corresponding localized low and high eigenmodes, $\lambda_{O}$ and $\tilde{\lambda}_{O}$, found in the presence of a single, isolated object; a clear anticorrelation is also seen near the mobility edges (see footnote 3). The position of the peaks in the correlation match almost perfectly those in the spectral density of low modes, as shown in Fig. 21 (second panel), while the matching seems less nice at the high end of the spectrum (Fig. 21, third panel). However, although relatively rare, the objects are not isolated from each other, and so the spectrum is not simply the union of the individual spectra of the configurations corresponding to the objects in Tab. 1, and the positions of the peaks in $\rho$ are expected to deviate from the isolated-object case. On the other hand, a comparison between the positions of the correlation peaks and the eigenvalues of the localized modes for isolated objects (fourth panel) shows that even in a heterogeneous environment the spectrum responds the most to the center-vortex configurations of Tab. 1 almost exactly where the “free” eigenvalue would be. The bigger deviation of the spectral density peaks from the correlation peaks for the high modes is then likely due to the fact that the localized high modes associated with isolated objects have eigenvalues closer to each other than the corresponding localized low modes, and therefore are more likely to mix when several objects are present, distorting the peak structure from the isolated-object case more than at the low end of the spectrum.
@@
Figure 22: Normalized spectral density at fixed number $\mathcal{N}$ of vertical squares (top panel), and its difference with the full normalized spectral density (bottom panel). The dashed line is at $\lambda\simeq 0.55$, which is the eigenvalue of the localized low mode corresponding to an isolated vertical square; the solid line indicates the position of the mobility edge. Here $\beta=0.74$, $N_{s}=20$, and $N_{t}=4$.
@@
Since $N_{O}V_{O}/V<1$ is the fraction of the lattice occupied by objects, the correlation is positive.
Interestingly, in Fig. 24 a positive correlation between low and high modes, particularly strong near $(\lambda_{1},\lambda_{2})=(\lambda_{O},\tilde{\lambda}_{O})$, is observed also in the confined phase. As a nonzero density of finite clusters is present also in the confined phase (see Sec. III.2), it is not surprising that they support the nonzero spectral density of localized high modes, as this result further confirms. It is somewhat surprising, though, that a strong correlation with the spectral regions where the low-mode counterparts would be found survives also in the confined phase, where low modes are delocalized. This indicates that low modes still carry some local structure highly correlated with the center-vortex configuration. This also explains their anticorrelation with the bulk modes.
@@
The results of this subsection show in further detail how the localization of low Dirac modes in the deconfined phase of $\mathbb{Z}_{2}$ gauge theory is explained by the opening of the spectral pseudogap due to Polyakov loop ordering, and by the presence of disconnected finite clusters, i.e., center vortices not spanning the whole lattice, that are the main localization centers. As both Polyakov-loop ordering and the fragmentation of the infinite center cluster are direct manifestations of deconfinement, the strong connection between deconfinement and localization is evident. Notice that center vortices being the main localization centers is not at all in contradiction with the sea/islands picture, as Polyakov loop, islands, and center vortices are very strongly correlated with each other, as shown in the previous subsection. Note, however, that the vortex structures associated with the lowest modes (see Fig. 20) correspond to more complex structures in the gauge configuration than a simple Polyakov-loop fluctuation. The simplest center vortex associated with such a fluctuation is the “horizontal square”, that is more relevant higher up in the spectrum. This shows that center vortices play a more fundamental role than Polyakov-loop fluctuations in setting up favorable localization centers for the low Dirac modes.
@@ IV Conclusions
In this paper we have investigated different aspects of the deconfinement transition in finite-temperature (2+1)-dimensional $\mathbb{Z}_{2}$ gauge theory on the lattice, and of its relation with the localization of low Dirac modes.
As a first step, we confirmed the presence of an Anderson transition in the spectrum of the staggered Dirac operator, separating localized low-lying modes from delocalized bulk modes, in the deconfined phase of the theory. The presence of localized low modes, appearing at the deconfinement transition, was shown in Ref. [19], short only of a finite-size scaling analysis. Here we performed this analysis, fully validating the claims of Ref. [19].
Our results are interesting also from the point of view of disordered systems, independently of their connection with the confinement of quarks. Our finite-size-scaling analysis shows that (2+1)-dimensional $\mathbb{Z}_{2}$ gauge theory is another (spatially) two-dimensional disordered system in the orthogonal class displaying a BKT-type Anderson transition [73, 75]. Furthermore, the dependence of the position of the mobility edge on the temperature is rather counterintuitive, and different from what has been observed so far in lattice gauge theories. One surprising aspect is that our results are consistent with the sudden appearance of the mobility edge at a finite distance from the origin of the spectrum at the critical temperature. Such a behavior seems natural for a first-order deconfinement transition [16, 17], but quite less so for the second-order one found in the present model.
We then further established the characterization of the deconfinement transition as a percolation transition of center vortices [42, 43, 54] (see also Ref. [57]), determining the percolation temperature and finding it in agreement with the deconfinement temperature [34]. An interesting finding is that the critical properties of the vortex percolation transition are different from those of ordinary two-dimensional percolation [60, 61, 62, 63]. This is probably due to the long-range nature of the constraint on the percolating objects, i.e., the requirement that active bonds form closed loops on the dual lattice.
@@
In (2+1)-dimensional $\mathbb{Z}_{2}$ gauge theory the very close connection between confinement and specific topological objects (center vortices, in this case) is almost trivially expected, due to the simplicity of the dynamics (it is nonetheless nice to be able to confirm it explicitly). This allowed us to investigate in an inequivocal way (albeit, perhaps, in too simplified an environment) the connection between deconfinement and localization of Dirac modes, strongly suggested by a wealth of results [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. Specifically, we looked directly at the correlation between localized modes and the center vortices whose percolation signals confinement. Such a correlation is demonstrated here at the microscopic level, identifying the structures in the gauge configurations mainly responsible for localizing the Dirac modes: these are precisely the most elementary center vortices. This is not trivial: localization could have been more directly related to the local fluctuations of the (temporal) Polyakov loops, as expected in the original sea/islands picture [5, 30, 31, 32, 24], which are correlated with center vortices but do not coincide with them. The resulting picture is then that as the temperature is lowered and the confinement transition is approached, such vortices approach a percolation transition, and the localized modes they support tend to mix more strongly, therefore tending to delocalize. The connection between the confinement-deconfinement transition and delocalization-localization transition is then fully elucidated by identifying the gauge structures that are responsible for both phenomena. This supports our hope that identifying the structures responsible for low-mode localization in more realistic gauge theories, including the physically relevant case of QCD, can lead to a better understanding of the microscopic mechanism behind confinement.
@@ Appendix A Van Hove singularities
Due to the ordering of the Polyakov loop, in the deconfined phase the bulk of the spectrum should resemble that of the free theory in the physical sector, and that of the free theory supplemented by a uniform nontrivial Polyakov loop in the unphysical sector. This includes reproducing the Van Hove singularities [80] of the spectral density where the “band energies” $\mathcal{E}_{n}(\vec{p}\,)=\sqrt{(\sin\omega_{n})^{2}+(\sin p_{1})^{2}+(\sin p_{2})^{2}}$ have a vanishing gradient, $\vec{\nabla}\mathcal{E}_{n}(\vec{p}\,)=[2\mathcal{E}_{n}(\vec{p}\,)]^{-1}(\sin 2p_{1},\sin 2p_{2})=0$. Here the various bands are characterized by $\omega_{n}^{\mathrm{phys}}=\frac{(2n+1)\pi}{N_{t}}$ for the physical sector, and $\omega_{n}^{\mathrm{unphys}}=\frac{2n\pi}{N_{t}}$ for the unphysical sector, with $n=0,\ldots,N_{t}-1$ in both cases. The spatial momentum $\vec{p}=(p_{1},p_{2})$ is restricted to the first Brillouin zone; in the large-volume limit, $p_{i}\in[-\pi,\pi]$.